# Universal behavior of the Covid-19 tails: Inverse power-law distribution

### v1

@@ Abstract
Power-law distribution is one of the most important laws known in nature. Such a special universal behavior is known to occur in very few physical systems. In this work, we analyzed the mortality distribution of the Covid-19 pandemic tails for different countries and continents to discuss the possible universal behavior of the pandemic. Surprisingly, we found that the mortality distribution of Covid-19 follows inverse power-law decays. These universal behaviors for the pandemic are reported in the present work for the first time. Additionally, we showed that mortality tails also decay with time obeying to the inverse power law.
@@ I Introduction
In this study, we reported that the inverse power law distribution can be seen in another phenomenon. As known, Covid-19, which started in 2020, was declared a pandemic by the World Health Organization [22]. Between 2020 and 2023, it can be said that it is recorded that approximately one billion people were infected and more than seven million people lost their lives [23]. We examined the tail of Covid-19, that is, a short region near the end of the Covid-19 pandemic, and we showed that the distribution function in this region complies with the inverse power-law. In addition, the time dependence of deaths occurring in this tail region will be examined in detail.
@@ II Power-law distribution of the Covid-19 tails
where $\alpha$ was obtained separately for each country. This universal behavior is known as the inverse power-law.
@@
As can be seen from the sub-figures, $\alpha$ values take values between 0 and 2 for countries. For instance; for DEU $\alpha=1.0347$, for ESP $\alpha=1.0347$, for FRA $\alpha=1.0347$, for GBR $\alpha=1.0347$, for GRC $\alpha=1.0347$, for ITA $\alpha=1.0347$, for NLD $\alpha=1.0347$, and for TUR $\alpha=1.0347$.
@@ III Time-dependent Decay of the Covid-19 tails
As in the previous section, we analyze eight same countries and the same continental data to discuss the time dependence of the Covid-19 tails. Therefore, we analyze the time dependence of the Covid-19 data tails for the same countries given in the previous section.
@@
It is important to see how the deaths resulting from the Covid-19 epidemic in these countries have decreased over time recently. Because we showed that the death distributions for this period obeyed the inverse power law. Therefore, the time-dependent decay of this period is a sign that it may comply with a universal law. Time dependence for these countries is given in Fig.3. These results show that the time-dependent decay of the Covid-19 tail obeys the power law:
@@
where $\beta$ is the decay exponent.
@@ IV Conclusion
In this study, it is the first time, we analyzed the tails of Covid-19 for eight countries and six continental regions to discuss the universal behavior in the distribution and time dependence of the tails. Firstly, we show that the deaths in the tail of Covid-19 for the different countries and different regions we analyzed can be represented by a distribution function that obeys the inverse power law: $P(m)\sim m^{-\alpha}$ with $20<\alpha<$ 2. Secondly, we analyzed the time dependence of the tail of Covid-19 death data for the same countries and the same continents and showed that the number of deaths decreases with an inverse power law: $M(t)\sim t^{-\beta}$ with $20<\beta<$ 2. These novel and interesting results are first reported in this work.
It should be pointed out that the inverse power law is universal, therefore the obtained results indicate that the tail of Covid-19 has a distribution function of the mortality and mortality relaxation with time to obey this universal law. On the other hand, we analyzed the mortality peaks in the entire time series as they increased and decreased, but we did not observe a similar universal behavior. The results we obtained are only valid in the Covid-19 tail, that is, in regions very close to the end of the pandemic.
@@
To the best of our knowledge, power law distribution and decay mechanisms are a controversial issue. Many alternative mechanisms have been suggested to generate power-law behavior. For example, it has been shown in various studies that the sum of different exponential distributions can generate a power-law distribution [24, 25], and that a random walk and critical phenomena [26] or taxonomy model of biological species [27, 28] can generate a power-law behavior. Additionally, it is possible to add different models [29, 30, 31] to them. We do not currently provide a mathematical model of how various stochastic effects produce power-laws in the tail of Covid-19 in our example, however, we can discuss possible predictions. The Covid-19 pandemic, which started in 2020 and infected approximately one billion people and caused more than seven million deaths by the end of 2023, has been thoroughly suppressed thanks to the preventive measures of governments and the developed vaccines. Despite the emergence of different variants, the virus rose to an acceptable level of herd immunity and lost its effect. The tail region, where Covid-19 begins to lose its effect, is therefore expected to differ significantly from the dynamics of the intermediate waves in the time series. Because we can evaluate that the fact that all stochastic parameters that play a role in suppressing the lethality of Covid-19 in the tail region have been maximized may have played a role in generating the power-law dynamics.

### v2

@@ Abstract
Power-law distribution is one of the most important laws known in nature. Such a special universal behavior is known to occur in very few physical systems. In this Letter, we analyzed the mortality distribution of the Covid-19 pandemic tails for different countries and continents to discuss the possible universal behavior of the pandemic. Surprisingly, we found that the mortality distribution of Covid-19 final i.e., the latest tails in 2023 follows inverse power-law decays. These universal behaviors for the pandemic are reported in the present work for the first time. Additionally, we showed that these mortality tails also decay with time obeying to the inverse power-law.
@@ I Introduction
In this Letter, we reported that the inverse power-law distribution can be seen in Covid-19 tails. As known, Covid-19, which started in 2020, was declared a pandemic by the World Health Organization [22]. Between 2020 and 2023, it can be said that it is recorded that approximately one billion people were infected and more than seven million people lost their lives [23]. We examined the final i.e., the latest tails in 2023 of Covid-19, that is, a short region near the end of the Covid-19 pandemic in 2023, and we showed that the distribution function in this region complies with the inverse power-law. In addition, the time dependence of deaths occurring in this tail region will be examined in detail.
@@ II Power-law distribution of the Covid-19 tails
where $m$ denotes mortality frequency, and $\alpha$ is the exponent that characterizes mortality rates for each country. This universal behavior is known as the inverse power-law.
@@
As can be seen from the sub-figures, $\alpha$ values take values between 0 and 2 for countries. For instance; for DEU $\alpha=1.4171$, for ESP $\alpha=1.543$, for FRA $\alpha=1.4721$, for GBR $\alpha=1.4424$, for GRC $\alpha=0.90999$, for ITA $\alpha=1.3557$, for NLD $\alpha=1.4359$, and for TUR $\alpha=0.77581$.
@@ III Time-dependent Decay of the Covid-19 tails
As in the previous section, we analyze eight same countries and the same continental data to discuss the time dependence of the Covid-19 tails. Therefore, we also analyze the time dependence of the Covid-19 data tails for the same countries given in the previous section.
@@
It is important to see how the deaths resulting from the Covid-19 epidemic in these countries have decreased over time recently. Because, we showed that the death distributions for this period obeyed the inverse power-law. Therefore, the time-dependent decay of this period is a sign that it may comply with a universal law. Time dependence for these countries is given in Fig.3. These results show that the time-dependent decay of the Covid-19 tail obeys the power-law:
@@
where $t$ is the re-scaled time step and $\beta$ is the decay exponent of the mortality for different countries.
@@ IV Discussion and Conclusion
In this study, we analyzed the last tails of Covid-19 for eight countries and six continental regions to discuss the universal behavior in the distribution and time dependence of the tails. Firstly, we show that the deaths in the latest tails of Covid-19 in 2023 for the different countries and different regions we analyzed can be represented by a distribution function that obeys the inverse power-law: $P(m)\sim m^{-\alpha}$ with $20<\alpha<$ 2. Secondly, we analyzed the time dependence of the tail of Covid-19 death data for the same countries and the same continents and showed that the number of deaths decreases with an inverse power-law: $M(t)\sim t^{-\beta}$ with $20<\beta<$ 2. These novel and interesting results are reported in this work.
It should be noted that, so far, many statistics have been applied to Covid-19 data to estimate the spreading dynamics of the pandemic. Weibull and modified Weibull distributions can be considered among the most widely used [24, 25, 26, 24, 27]. Similarly, different distributions were applied to the epidemic data. For example; Kaniadakis et al [28] apply $\kappa$-statistics distribution to some empiric Covid-19 data from several countries. They showed that empirical data of Covid-19 peaks are consistent with $\kappa$-statistics distribution. Similarly, Tsallis and Tirnakli [29] proposed a $q$-statistical functional form to describe Covid-19 data for several countries [29]. In a recent study, Kaniadakis also showed that when power-law interactions are taken into account, the SIR model can give more harmonious results and the compartment population distributions evolve in time exhibiting power-law tails [30]. On the other hand, in many studies in the literature, it has been reported that some of the growth or decay tails of the independent waves of the epidemic data between 2020 and 2023 are exponential [31, 32, 33, 34], sub-exponential [35] or some follow the inverse power-law decay [31, 36, 28, 29, 30]. These analyses may indicate that Covid-19 data of the different countries may have different characteristics at the same periods. Unlike previous studies, in this study, we were only interested in the final Covid-19 peak tails in 2023 for various countries and continentals not independent Covid-19 peaks between 2020 and 2023 periods, and we report that the latest tails of Covid-19 data decay with power-law like form which may denotes universal character as well in the previous studies [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 8, 11, 12, 13].
@@
If we return to our work again, to the best of our knowledge, we can state that the physical origin of the power law distribution and decay mechanisms are a controversial issue. Many alternative mechanisms have been suggested to generate power-law behavior so far. For example, it has been shown in various studies that the sum of different exponential distributions can generate a power-law distribution [37, 38], and that a random walk and critical phenomena [39] or taxonomy model of biological species [40, 41] can generate a power-law behavior. Additionally, it is possible to add different models [42, 43, 44] to them. In this work, we do not currently provide a mathematical model of how various stochastic effects produce power-laws in the tail of Covid-19 in our example, however, we can discuss possible predictions. The Covid-19 pandemic, which started in 2020 and infected approximately one billion people and caused more than seven million deaths by the end of 2023, has been thoroughly suppressed thanks to the preventive measures of governments and the developed vaccines. Despite the emergence of different variants, the virus rose to an acceptable level of herd immunity and lost its effect. The tail region, where Covid-19 begins to lose its effect, is therefore expected to differ significantly from the dynamics of the intermediate waves in the time series. Because, we can evaluate that the fact that all stochastic parameters that play a role in suppressing the lethality of Covid-19 in the tail region have been maximized may have played a role in generating the power-law dynamics.